# Multi-Order Wavelet Derivative Transform for Deep Time Series Forecasting

### v1

@@ Abstract
In deep time series forecasting, the Fourier Transform (FT) is extensively employed for frequency representation learning. However, it often struggles in capturing multi-scale, time-sensitive patterns. Although the Wavelet Transform (WT) can capture these patterns through frequency decomposition, its coefficients are insensitive to change points in time series, leading to suboptimal modeling. To mitigate these limitations, we introduce the multi-order Wavelet Derivative Transform (WDT) grounded in the WT, enabling the extraction of time-aware patterns spanning both the overall trend and subtle fluctuations. Compared with the standard FT and WT, which model the raw series, the WDT operates on the derivative of the series, selectively magnifying rate-of-change cues and exposing abrupt regime shifts that are particularly informative for time series modeling. Practically, we embed the WDT into a multi-branch framework named WaveTS, which decomposes the input series into multi-scale time-frequency coefficients, refines them via linear layers, and reconstructs them into the time domain via the inverse WDT. Extensive experiments on ten benchmark datasets demonstrate that WaveTS achieves state-of-the-art forecasting accuracy while retaining high computational efficiency.
@@ 1 Introduction
Time series forecasting aims to predict future values by leveraging patterns of past observations, playing a pivotal role in diverse fields such as climate, finance, and transportation [42, 60, 44, 50, 37, 18]. In recent years, deep learning techniques have gained prominence for modeling long-term temporal dependencies, reflected in MLP-based [58, 59], RNN-based [16, 12], Transformer-based [26, 32, 48] etc., underscoring the effectiveness of representation learning in the time domain.
Figure 1: (a)–(d) Distinct temporal variations may yield identical frequency spectra after applying the Fourier transform, revealing the spectral ambiguity. (e)(f) The Fourier-derivative makes the spectrum stationary yet discards macro-trend information. (e)(g)(h) The Wavelet Derivative Transform (WDT) retains those trends while offering complementary detail. (i) Relative to wavelet coefficients in the standard wavelet transform (blue curve), WDT coefficients pinpoint abrupt regime shifts (red curve), yielding a sharper cue for non-stationary dynamics.
@@
Alongside these time-domain methods, recent deep forecasting models have also been proposed to model time series in the frequency domain, namely Frequency domain Representation Learning (FRL). These FRL-based forecasters largely leverage the Fourier transform for domain transformation, thereby extracting the frequency components. Subsequently, they process these frequency components using complex-valued neural networks to derive semantic frequency representations. For example, FreTS [56] introduces frequency domain MLPs to effectively model both temporal and inter-variable dependencies. FITS [49] incorporates frequency interpolation through a complex-valued linear layer, rendering it 50 times more compact than previous linear models. Additionally, several studies have incorporated the FRL paradigm into more complex architectures like Transformer and Graph Neural Network to extract more semantic and robust features in the frequency domain [63, 6, 3, 55].
@@
However, previous FRL-based forecasters exhibit a heavy reliance on the Fourier transform for domain transformation [56, 49, 45, 53, 7, 8]. While this methodology is both effective and computationally efficient, it is inherently constrained in its ability to model time-sensitive patterns. As illustrated in Figure 1(a)-(d), similar frequency spectra are generated after the application of the Fourier transform to time series exhibiting markedly distinct temporal variations. This leads to ambiguities in deep representation learning, as well as inadequate modeling of the time-frequency characteristics and non-stationarity of time series. DeRiTS [7] attempts to address this issue by proposing a deep Fourier derivative learning framework aimed at learning stationary frequency representations (Figure 1(f)). However, this approach neglects to consider macro temporal patterns, such as the overall trend, and still fails to capture time-aware features, ultimately resulting in inaccurate forecasting.
@@
To address the aforementioned issues, this paper proposes a new transformation for time series, namely Wavelet Derivative Transform (WDT). WDT is built upon the Wavelet Transform (WT), targeted in modeling the derivative of time series rather than the raw series. It decomposes time series into slowly varying trend coefficients and stationary oscillatory coefficients, revealing hierarchical (macro) and stationary (detail) patterns (Figure 1 (g) and (h)). Moreover, as shown in Figure 1 (i), the coefficient of standard WT (blue curve) predominantly smooths the signal and simply fits its overall trend, ignoring the depiction of abrupt change points in black dotted boxes. However, by encoding the series’ derivative, the coefficient of WDT (red curve) not only preserves the macroscopic trend, but also highlights local bursts more sharply than the standard wavelet coefficient, providing richer cues for pinpointing sharp discontinuities, thereby facilitating its perception of local dynamics. Based on WDT, we propose a multi-branch architecture WaveTS, where each branch corresponds to a WaveTS block designated to process a specific order of WDT. WDT firstly decomposes time series into time-frequency coefficients. Subsequently, Frequency Refinement Units, comprising multiple hardware-friendly real-valued linear layers, refine these coefficients into more granular time-frequency representations, enabling WaveTS to extract richer patterns across various spectral bands. Ultimately, the inverse WDT reconstructs both historical and future sequences back into the time domain. Notably, the reversible nature of the WDT guarantees that no energy and information is lost, enhancing effective and efficient modeling. Our contributions can be articulated as follows:
@@
  • We propose a multi-order Wavelet Derivative Transform that captures multi-scale time-frequency patterns, enabling sharper localization of regime shifts than the standard wavelet transform.
@@ 2.1 Wavelet Transform
The Discrete Wavelet Transform (DWT) [29] decomposes a discrete time series signal into time-frequency components. Given a time series $[X(1),\dots,X(N)]$ of length $N$, the DWT coefficient $W_{k,j}$ at scale $k$ and translation $j$ is given by:
@@
where $\psi_{k,j}[n]=2^{-k/2}\,\psi\!\bigl{[}2^{-k}n-j\bigr{]}$ is the discrete basis function with mother wavelet $\psi[n]$, and $K=\lfloor\log_{2}N\rfloor$ is the maximum dyadic scale. At each scale $k$, the signal is split into a high-frequency (detail) component capturing the abrupt changes, and a low-frequency (approximation) component reflecting the overall trends. Multi-resolution analysis of both the time and frequency components is obtained by repeatedly decomposing the low-frequency components. The Haar [11] and Daubechies [4] wavelets are commonly used. They provide orthogonal or nearly orthogonal decompositions, ensuring that the DWT coefficients at different scales and translations capture non-redundant features of the signal [23].
2.2 Related Works
@@
Deep learning-based methodologies have been extensively applied to time series forecasting, as evidenced by TCNs [47, 24], and Transformer-based models [26, 32, 43, 61]. More recently, LLM-based methods [64, 17, 21] have emerged, leveraging the capacity of large pre-trained language models for knowledge transfer. Alongside these, FRL-based approaches [56, 49, 55, 53, 54, 62] employ the Fourier transform to uncover frequency patterns implicit in time-domain signals. FAN [52] identifies non-stationary factors by extracting dominant real frequency components, while DeRiTS [7] represents time series through frequency derivation However, existing FRL-based models still face challenges in capturing the inherently multi-scale time-frequency features, leading to ineffective time series modeling.
@@
In contrast to Fourier-based transformations, the wavelet transform decomposes signals into localized time-frequency components, making it well-suited for multi-scale analysis [63, 38, 9, 34]. For example, AdaWaveNet [57] employs an adaptive and learnable wavelet decomposition scheme to address non-stationary data, while T-WaveNet [30] adopts a tree-structured network for iteratively disentangling dominant frequency subbands. WPMixer [31] couples multi-resolution wavelet decomposition with patching and MLP mixing. wMDN [40] embeds wavelet-based frequency analysis into a deep learning framework through fine-tuning.
@@ 3 Methodology
Though standard wavelet-based forecasters can provide a hierarchical scale-by-scale decomposition, their coefficients mainly encode amplitude and thus under-emphasize multi-scale rates of change. This makes them less effective at highlighting local regime shifts and at modeling non-stationary temporal patterns or time-varying distribution shift within a time series (i.e., the change of its mean, variance over time). In this paper, we propose the Wavelet Derivative Transform (WDT) (section 3.1), which explicitly models the derivatives of a time series to sharpen multi-scale change signals. Building on it, we further propose WaveTS, a multi-branch architecture built upon WaveTS blocks (section 3.2.2) that applies multi-order WDT, which linearly refines the resulting coefficients through Frequency Refinement Unit, and reconstructs them back to the time domain for accurate and efficient backcasting while forecasting (section 3.2.3).
In the following, we denote a multivariate time series by $[X_{1},X_{2},\ldots,X_{T}]\in\mathbb{R}^{C\times T}$, where each $X_{t}\in\mathbb{R}^{C}$ corresponds to the observations of $C$ variates at the $t$th time step. We use a backcasting and forecasting approach to align the model forecasts with the observed historical trend [3]. Specifically, at time $t$, the model input $\mathbf{X}_{t}=[X_{t-L+1},X_{t-L+2},\ldots,X_{t}]\in\mathbb{R}^{C\times L}$ is a window of $L$ observations, and the model output is $\mathbf{Y}_{t}=[\hat{X}_{t-L+1},\dots,\hat{X}_{t-1},\hat{X}_{t},\hat{X}_{t+1},\ldots,\hat{X}_{t+\tau}]\in\mathbb{R}^{C\times L+\tau}$, where $[\hat{X}_{t-L+1},\dots,\hat{X}_{t-1},\hat{X}_{t}]$ are the backcasting results and $[\hat{X}_{t+1},\ldots,\hat{X}_{t+\tau}]$ are the forecasting results. The model, parameterized by $\theta$, denoted $f_{\theta}(\cdot)$, utilizes the historical data $\mathbf{X}_{t}$ to estimate the backcasting and forecasting values $\mathbf{Y}_{t}=f_{\theta}(\mathbf{X}_{t})$.
@@ 3.1 Wavelet Derivative Transform (WDT)
However, computing the derivative directly in the time domain is fragile: finite derivatives amplify high-frequency noise, shorten the usable sequence and complicate boundary handling. WDT avoids these pitfalls by performing derivative of the wavelet basis instead of the raw series—a spectral-domain trick that keeps the transform perfectly invertible, suppresses numerical noise and preserves scale alignment. Put simply, the WT of the multi-order derivative of $X(t)$, can be expressed as the WT of $X(t)$ using the multi-order derivative of the mother wavelet. Therefore, WDT captures derivative information without the usual time-domain baggage. See detailed proof in Appendix A.1.
Definition 3.1 (Wavelet Derivative Transform (WDT)).
@@
[…]
Definition 3.2 (Inverse Wavelet Derivative Transform (iWDT)).
@@
[…]
where $\mathcal{D}^{(n)}(W_{k,j})$ denotes the $n$-th-order WDT coefficient at scale $k$ (there are total $K$ scales) and translation $j$, $\psi_{k,j}(t)=2^{-k/2}\psi(2^{-k}t-j)$ is the standard discrete wavelet basis associated with the mother wavelet $\psi$, and $\tilde{W}^{(n)}_{k,j}\!=\!(-1)^{-n}2^{-nk}\mathcal{D}^{(n)}(W_{k,j})$ is the rescaled coefficient that matches a conventional inverse WT. Because the rescaling is exact and the inverse WT is perfectly reconstructive, iWDT restores $X(t)$ without information loss; a Parseval-type proof of its energy-conservation property is provided in Appendix A.2.

### v2

@@ Abstract
In deep time series forecasting, the Fourier Transform (FT) is extensively employed for frequency representation learning. However, it often struggles in capturing multi-scale, time-sensitive patterns. Although the Wavelet Transform (WT) can capture these patterns through frequency decomposition, its coefficients are insensitive to abrupt changes in the time series, leading to suboptimal modeling. To mitigate these limitations, we introduce the multi-order Wavelet Derivative Transform (WDT) grounded in the WT, enabling the extraction of time-aware patterns spanning both the overall trend and subtle fluctuations. Compared with the standard FT and WT, which model the raw series, WDT operates on the derivative of the series, selectively magnifying rate-of-change cues and exposing abrupt regime shifts that are particularly informative for time series modeling. Practically, we embed the WDT into a multi-branch framework named WaveTS, which decomposes the input series into multi-scale time-frequency coefficients, refines them via linear layers, and reconstructs them into the time domain via the inverse WDT. Extensive experiments on multiple benchmark datasets demonstrate that WaveTS achieves state-of-the-art forecasting accuracy while retaining high computational efficiency. The code is available at https://github.com/zhouziyu02/WaveTS.
@@ 1 Introduction
Time series forecasting aims to predict future values by leveraging patterns of past observations, playing a pivotal role in diverse fields such as climate, finance, and transportation (Wang et al., 2020; Zhang et al., 2023; Wang et al., 2024c; Yang et al., 2025; Qiu et al., 2025; Kieu et al., 2019). In recent years, deep learning techniques have gained prominence for modeling long-term temporal dependencies, reflected in MLP-based (Zeng et al., 2023; Zhang et al., 2022), RNN-based (Jia et al., 2023; Hewamalage et al., 2021), Transformer-based (Liu et al., 2024c; Nie et al., 2023; Wu et al., 2024) etc., underscoring the effectiveness of temporal representation learning in the time domain.
@@
Alongside these time-domain methods, recent deep forecasting models have also been proposed to model time series in the frequency domain, namely Frequency domain Representation Learning (FRL). These FRL-based forecasters largely leverage the Fourier transform for domain transformation, thereby extracting the frequency components. Subsequently, they process these frequency components using complex-valued neural networks to derive semantic frequency representations. For example, FreTS (Yi et al., 2023b) introduces frequency domain MLPs to effectively model both intra-variate and inter-variate dependencies. FITS (Xu et al., 2024) incorporates frequency interpolation through a complex-valued linear layer, rendering it 50 times more compact than previous linear models. Additionally, several studies have incorporated the FRL paradigm into more complex architectures like Transformers and Graph Neural Networks to extract more semantic and robust features in the frequency domain (Zhou et al., 2022b; Eldele et al., 2024; Cao et al., 2020; Yi et al., 2023a).
@@
However, previous FRL-based forecasters exhibit a heavy reliance on the Fourier transform for domain transformation (Yi et al., 2023b; Xu et al., 2024; Wang et al., 2025c; Yi et al., 2024a; Fan et al., 2014; Fei et al., 2025). While this methodology is both effective and computationally efficient, it is inherently constrained in its ability to model time-sensitive patterns. As illustrated in Figure 1 (a)-(d), similar frequency spectra are generated after the application of the Fourier transform to time series exhibiting markedly distinct temporal variations. This leads to ambiguities in deep representation learning, as well as inadequate modeling of the time-frequency characteristics and non-stationarity of time series. DeRiTS (Fan et al., 2014) attempts to address this issue by proposing a deep Fourier derivative learning framework aimed at learning stationary frequency representations (Figure 1 (f)). However, this approach neglects to consider macro temporal patterns, such as the overall trend, and still fails to capture time-aware features, ultimately resulting in inaccurate forecasting.
To address the aforementioned issues, this paper proposes a new transformation for time series, namely Wavelet Derivative Transform (WDT). WDT is built upon the Wavelet Transform (WT), targeted in modeling the derivative of time series rather than the raw series. It decomposes time series into slowly varying trend coefficients and stationary oscillatory coefficients, revealing hierarchical (macro) and stationary (detail) patterns (Figure 1 (g) and (h)).
@@
Moreover, as shown in Figure 1 (i), the coefficient of standard WT (blue curve) predominantly smooths the signal and simply fits its overall trend, ignoring the depiction of abrupt change points in black dotted boxes. However, by encoding the series’ derivative, the coefficient of WDT (red curve) not only preserves the macroscopic trend, but also highlights local bursts more sharply than the standard wavelet coefficient, providing richer cues for pinpointing sharp discontinuities, thereby facilitating its perception of local dynamics. Based on WDT, we propose a multi-branch architecture WaveTS, where each branch corresponds to a WaveTS block designated to process a specific order of WDT. WDT first decomposes time series into time-frequency coefficients. Subsequently, Frequency Refinement Units, comprising multiple hardware-friendly real-valued linear layers, refine these coefficients into more granular time-frequency representations, enabling WaveTS to extract richer patterns across various spectral bands. Ultimately, the inverse WDT reconstructs both historical and future sequences back into the time domain. To make this multi-order construction mathematically sound, we rigorously analyze WDT and prove that it defines a well-posed linear transform with an exact inverse and stable multi-order coefficients. These results clarify when WDT preserves the energy and information content of the original series and explain why its derivative-aware coefficients can be safely exploited in downstream models. Our contributions can be articulated as follows:
Figure 1: (a)–(d) Distinct temporal variations may yield identical frequency spectra after applying the Fourier transform, revealing spectral ambiguity; (e) original time series; (f) the Fourier Derived Series only exhibits high-frequency fluctuations with values around 0 and fails to capture macro temporal patterns such as the overall trend of the original series in (e); (g) the Wavelet Derivative Transform extracts macro low-frequency trend patterns; (h) the Wavelet Derivative Transform extracts micro high-frequency fluctuations; (i) relative to the wavelet coefficients in the standard wavelet transform (blue curve), WDT coefficients (red curve) pinpoint abrupt regime shifts, yielding a sharper cue for non-stationary dynamics.
@@
  • We propose a multi-order Wavelet Derivative Transform that captures multi-scale time–frequency patterns, enabling sharper localization of regime shifts than the standard wavelet transform.
  • We provide a rigorous theoretical analysis of the proposed Wavelet Derivative Transform, proving the existence of an exact inverse transform and establishing energy-preserving properties, thereby clarifying the mathematical difficulty and significance of our construction.
@@ 2.1 Wavelet Transform
The Discrete Wavelet Transform (DWT) Mallat (1989) decomposes a discrete time series signal into time-frequency components. Given a time series $X(1),\dots,X(N)]$ of length $N$, the DWT coefficient $W_{k,j}$ at scale $k$ and translation $j$ is given by:
@@
where $\psi_{k,j}[n]=2^{-k/2}\,\psi\bigl[2^{-k}n-j\bigr]$ is the discrete basis function with mother wavelet $\psi[n]$, and $K=\lfloor\log_{2}N\rfloor$ is the maximum dyadic scale. At each scale $k$, the signal is split into a high-frequency (detail) component capturing the abrupt changes, and a low-frequency (approximation) component reflecting the overall trends. Multi-resolution analysis of both the time and frequency components is obtained by repeatedly decomposing the low-frequency components. The Haar (Haar, 1911) and Daubechies (Daubechies, 1988) wavelets are commonly used. They provide orthogonal or nearly orthogonal decompositions, ensuring that the DWT coefficients at different scales and translations capture non-redundant features of the signal (Liu et al., 2024a).
2.2 Related Work
@@
Deep learning-based methodologies have been extensively applied to time series forecasting, as evidenced by TCNs (Wu et al., 2023; Liu et al., 2022), and Transformer-based models (Liu et al., 2024c; Nie et al., 2023; Wang et al., 2024b; Zhou et al., 2021). More recently, LLM-based methods (Zhou et al., 2023; Jin et al., 2024; Liang et al., 2024) have emerged, leveraging the capacity of large pre-trained language models for knowledge transfer. Alongside these, FRL-based approaches (Yi et al., 2023b; Xu et al., 2024; Yi et al., 2023a; 2024a; 2024b; Zhou et al., 2022a; 2026) employ the Fourier transform to uncover frequency patterns implicit in time-domain signals. FAN (Ye et al., 2024) identifies non-stationary factors by extracting dominant real frequency components, while DeRiTS (Fan et al., 2014) represents time series through frequency derivation. However, existing FRL-based models still face challenges in capturing the inherently multi-scale time-frequency features, leading to ineffective time series modeling.
@@
In contrast to Fourier-based transformations, the wavelet transform decomposes signals into localized time-frequency components, making it well-suited for multi-scale analysis (Zhou et al., 2022b; Stefenon et al., 2023; Gao et al., 2021; Ouyang et al., 2021). For example, wMDN (Wang et al., 2018) embeds wavelet-based frequency analysis into a deep learning framework through fine-tuning. AdaWaveNet (Yu et al., 2024) employs an adaptive and learnable wavelet decomposition scheme to address non-stationary data, while T-WaveNet (Liu et al., 2021) adopts a tree-structured network for iteratively disentangling dominant frequency subbands. WPMixer (Murad et al., 2025) couples multi-resolution wavelet decomposition with patching and MLP mixing. WaveletMixer (Zhang et al., 2025) is an iterative multi-level, multi-resolution, and multi-phase design that captures long- and short-term dependencies. However, most wavelet-coefficient pipelines do not explicitly address regime shifts and thus perform worse around transitions, even if overall accuracy is competitive. Augmenting them with shift-sensitive mechanisms (e.g., derivative-enhanced transforms) improves alignment near abrupt changes.
@@ 3 Methodology
Though standard wavelet-based forecasters can provide a hierarchical scale-by-scale decomposition, their coefficients mainly encode amplitude and thus under-emphasize multi-scale rates of change. This makes them less effective at highlighting local regime shifts and at modeling non-stationary temporal patterns or time-varying distribution shift within a time series (i.e., the change of its mean, variance over time). In this paper, we propose the Wavelet Derivative Transform (WDT) in Section 3.1 with its inverse, which explicitly models the derivatives of a time series to sharpen multi-scale change signals. Building on it, we further propose WaveTS, a multi-branch architecture built upon WaveTS blocks (Section 3.2.2) that applies multi-order WDT, which linearly refines the resulting coefficients through Frequency Refinement Unit, and reconstructs them back to the time domain for accurate and efficient backcasting while forecasting (Section 3.2.3).
In the following, we denote a multivariate time series by $X_{1},X_{2},\ldots,X_{T}]\in\mathbb{R}^{T\times C}$, where each $X_{t}\in\mathbb{R}^{C}$ corresponds to the observations of $C$ variates at the $t$-th time step. We use a backcasting and forecasting approach to align the model forecasts with the observed historical trend (Cao et al., 2020). Specifically, at time $t$, the model input $\mathbf{X}_{t}=[X_{t-L+1},X_{t-L+2},\ldots,X_{t}]\in\mathbb{R}^{L\times C}$ is a window of $L$ observations, and the model output is $\mathbf{Y}_{t}=[\hat{X}_{t-L+1},\dots,\hat{X}_{t-1},\hat{X}_{t},\hat{X}_{t+1},\ldots,\hat{X}_{t+\tau}]\in\mathbb{R}^{(L+\tau)\times C}$, where $\hat{X}_{t-L+1},\dots,\hat{X}_{t-1},\hat{X}_{t}]$ are the backcasting results and $\hat{X}_{t+1},\ldots,\hat{X}_{t+\tau}]$ are the forecasting results. The model, parameterized by $\theta$, denoted $f_{\theta}(\cdot)$, utilizes the historical data $\mathbf{X}_{t}$ to estimate the backcasting and forecasting values $\mathbf{Y}_{t}=f_{\theta}(\mathbf{X}_{t})$.
@@ 3.1 Wavelet Derivative Transform (WDT)
However, computing the derivative directly in the time domain is fragile: finite derivatives amplify high-frequency noise, shorten the usable sequence and complicate boundary handling. WDT avoids these pitfalls by performing derivative of the wavelet basis instead of the raw series—a spectral-domain trick that keeps the transform perfectly invertible, suppresses numerical noise and preserves scale alignment. Put simply, the WT of the multi-order derivative of $X(t)$, can be expressed as the WT of $X(t)$ using the multi-order derivative of the mother wavelet. Therefore, WDT captures derivative information without the usual time-domain baggage. See detailed proof in Appendix B.1.
Definition 1 (Wavelet Derivative Transform (WDT)).
@@
$\mathrm{WDT}^{(n)}_{k,j}\;=\;(-1)^{\,n}\,2^{nk}\,\tilde{W}^{(n)}_{k,j}$.
Here, $\tilde{W}^{(n)}_{k,j}\;=\;\sum_{t=0}^{T-1}X(t)\,\psi^{(n)}_{k,j}(t)$ is the wavelet transform of $X(t)$ using the $n$-order derivative of the mother wavelet $\psi^{(n)}(t)$, where $\psi^{(n)}_{k}(t)$ is the $n$-order derivative of the discrete wavelet basis function, defined as: $\psi^{(n)}_{k,j}(t)\;=\;2^{\,k/2}\,\psi^{(n)}\!\bigl(2^{\,k}t-j\bigr).$
Definition 2 (Inverse Wavelet Derivative Transform (iWDT)).
@@
$X(t)=\mathrm{iWDT}_{K}^{(n)}\bigl(\mathcal{D}^{(n)}(W_{k,j})\bigr)=\sum_{k=1}^{K}\sum_{j}\frac{\mathcal{D}^{(n)}(W_{k,j})}{(-1)^{\,n}\,2^{nk}}\,\psi_{k,j}(t)=\sum_{k=1}^{K}\sum_{j}\tilde{W}^{(n)}_{k,j}\,\psi_{k,j}(t),$ (2)
where $\mathcal{D}^{(n)}(W_{k,j})$ denotes the $n$-th-order WDT coefficient at scale $k$ (there are total $K$ scales) and translation $j$, $\psi_{k,j}(t)=2^{-k/2}\psi(2^{-k}t-j)$ is the standard discrete wavelet basis associated with the mother wavelet $\psi$, and $\tilde{W}^{(n)}_{k,j}\!=\!(-1)^{-n}2^{-nk}\mathcal{D}^{(n)}(W_{k,j})$ is the rescaled coefficient that matches a conventional inverse WT. Because the rescaling is exact and the inverse WT is perfectly reconstructive, iWDT restores $X(t)$ without information loss; a Parseval-type proof of its energy-conservation property is provided in Appendix B.2.